OpenStudy (anonymous):

An automobile starter motor is connected to a 12.0 V battery. When the starter is activated it draws 150 A of current, and the battery voltage drops to 7.0 V. What is the battery’s internal resistance?

OpenStudy (anonymous):

@robtobey

OpenStudy (kropot72):

The voltage drop across the internal resistance is 12 - 7 = 5 volts. Internal resistance is given by: \[\Large R _{internal}=\frac{V}{I}=\frac{5}{150}=(you\ can\ calculate)\ \Omega\]